A collection of nickles, dimes and quarters contains three times as many nickles as dimes and four more quarters than nickles. The total value of the collection $9.00 How many quarters are there? Please let me see the equation that you arrived at first?
Answer by Paul(988)

5(3x)+10(x)+25(4+3x)=900
15x+10x+100+75x=900
100x+100=900
100x=800
x=8
Hence there are 8 dimes, just plug in the value for x and solve the rest.